OpenStudy (anonymous):

A point moves along a circle with speed,p=dm,{d=0.5 m/s^2}...p,d,m are just variables!! Find the total speed and acceleration of point at moment when it covers (1/10)th of the circle after beginning of motion..!!!(Non-Uniform Motion)

OpenStudy (anonymous):

@april115 @andriod09 @AnElephant @ayushkr2611 @blueocean18 @countonme123 @Coolsector @dan815 @e.cociuba @Falco276 @Fifciol @GoldEverything @GOODMAN @HaperFink22 @ilfy214 @ivancsc1996 @Luigi0210 @Loser66 @lncognlto

OpenStudy (ivancsc1996):

If it is a uniform circular motion then d. We assume p is a speed and m is time

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Then???

OpenStudy (ivancsc1996):

The answer is d since the acceleration is constant. Maybe you are thinking about the total speed?

OpenStudy (anonymous):

O_O....Oopsy!!

OpenStudy (anonymous):

Yes speed

OpenStudy (ivancsc1996):

Then, we can use this equation: \[\theta _{f}=\theta _{i}+\omega _{i} t+\frac{ 1 }{ 2 } \alpha t ^{2}\]Where theta is the angle, omega is the angular velocity and alpha is the angular acceleration. Since initial angle and the initial angular velocity are zero, when we clear t we get:\[\sqrt{\frac{ 2 \theta _{f} }{ \alpha }}=t\] Now, alpha is related to d (the linear acceleration) by:\[\alpha=\frac{ d }{ r }\]That means that you need the radius of the circle. Do you have it?

OpenStudy (anonymous):

r??

OpenStudy (anonymous):

t,then will equal \[2\sqrt{{\pi*r}/5}\]

OpenStudy (anonymous):

@ivancsc1996

OpenStudy (anonymous):

so v wil be \[\sqrt{{(pi*r)}/5}\]

OpenStudy (ivancsc1996):

No, \[m=\sqrt{\frac{ 2 \times \frac{ 2 \pi }{ 10 }}{ \frac{ d }{ r } }}=\sqrt{\frac{ 2 \pi r}{ 5d }}\]Then you substitute that back into the equation for p and you get:\[p=d \sqrt{\frac{ 2 \pi r }{ 5d }}=\sqrt{\frac{ 2 d \pi r }{ 5 }}\]

OpenStudy (anonymous):

It shud be 4*2x/10

OpenStudy (anonymous):

@april115 @andriod09 @AnElephant @ayushkr2611 @blueocean18 @countonme123 @Coolsector @dan815 @e.cociuba @Falco276 @Fifciol @GoldEverything @GOODMAN @HaperFink22 @ilfy214 @ivancsc1996 @Luigi0210 @Loser66 @lncognlto

OpenStudy (anonymous):

@androidonyourface @ash2326 @amistre64 @surjithayer @sorour

OpenStudy (anonymous):

@TuringTest